4. The quantity of coffee demanded, QD, depends on the price of coffee, Pe, and the price of tea, PT. The quantity of coffee supplied, 05, depends on the price of coffee, Pe, and the price of electricity, Pt, according to the following equation: O- 17-2 Pe 10 Pr a If the price of tea is $1.00 and the price of elecricity is S0.50, what is the equilibrium price and quantity of coffee? b. What is/are the endogenous variable(s) in this model? c. What is/are the exogenous variable(s) in this model?
Answer #1

(a) Plugging in given values,

QD = 17 - 2Pc + 10 x 1 = 17 - 2Pc + 10 = 27 - 2Pc

QS = 2 + 3Pc - 5 x 0.5 = 2 + 3Pc - 2.5 = 3Pc - 0.5

In equilibrium, QD = QS.

27 - 2Pc = 3Pc - 0.5

5Pc = 27.5

Pc = $5.5

Q = (3 x 5.5) - 0.5 = 16.5 - 0.5 = 16

(b) Endogenous variables are those variables which depend on other variables in the system. Here, price of coffee (Pc) and quantity of coffee (Qc) are the endogenous variables.

(c) Exogenous variables are those variables which does not depend on other variables in the system. Here, price of tea (PT) and price of electricity (PE) are the exogenous variables.